OpenStudy (anonymous):

WILL GIVE MEDALS . (: What is the quotient: (3x2 + 4x - 15) ÷ (x + 3) ? 3x + 5 3x - 5 3x + 1 3x - 1, r = 1

OpenStudy (anonymous):

\[\frac{3x^2 + 4x - 15}{x + 3}=\frac{3x^2 + 9x -5x- 15}{x + 3}\]\[=\frac{3x(x + 3) -5(x+3)}{x + 3}\]\[=\frac{(x + 3)(3x -5)}{x + 3}\]\[=3x -5\]

OpenStudy (anonymous):

It's 3x-5 because =2x^2+4x-15 =2x^2+9x-5x-15 =3x(x+3) - 5(x+3) =(3x-5)(x+3) / (x+3) =3x-5

OpenStudy (anonymous):

Sorry, it's 3x^2 for the two first solutions... made an error